# Operator World Models for Reinforcement Learning

### v1

@@ 5 Theoretical Analysis
We now show that POWR converges to the global maximizer of the RL problem in 1. To this end, we first identify a family of function spaces guaranteed to be compatible with the policies generated by Alg. 1. Then, we provide an extension of the result in [12] for inexact PMD to infinite state spaces $\mathcal{X}$, showing the impact of the action-value approximation error on the convergence rates. By leveraging the simulation lemma, we will relate this error for the case of the estimator introduced Prop. 3 to the approximation errors of ${T}_{n}$ and $r_{n}$. Finally, we will use recent advancements in the characterization of CMEs’ fast learning rates to bound the sample complexity of these latter quantities, yielding error bounds for POWR.
@@
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact set and let ${\mathcal{H}}=W^{2,s}(\mathcal{X})$ be the Sobolev space of smoothness $s>0$ (see e.g. [33]). Let $\pi_{t}(a|\cdot)$ and $\hat{q}_{\pi_{t}}(\cdot,a)$ belong to ${\mathcal{H}}$ for any $a\in\mathcal{A}$ and $\pi_{t}(a|x)>0$ for any $x\in\mathcal{X}$. Then the policy $\pi_{t+1}$ solution to the PMD update in 13 belongs to ${\mathcal{H}}$.
According to Thm. 5, Sobolev spaces offer a viable choice for compatibility with PMD-generated policies. This observation is further supported by the fact that Sobolev spaces of smoothness $s>d/2$ are so-called reproducing kernel Hilbert spaces (rkhs) (see e.g. [34, Ch. 10]). We recall that rkhs are always naturally equipped with a $\phi:\mathcal{X}\to{\mathcal{H}}$ such that the inner product $\left\langle{\phi(x)},{\phi(x^{\prime})}\right\rangle=k(x,x^{\prime})$ defines a so-called reproducing kernel, namely a positive definite function that is (usually) efficient to evaluate, even if $\phi(x)$ is high or infinite dimensional. For example, ${\mathcal{H}}=W^{2,s}(\mathcal{X})$ with $s=\lceil\frac{d}{2}\rceil$ has associated kernel $k(x,x^{\prime})=e^{-\|x-x^{\prime}\|/\sigma}$ with bandwidth $\sigma>0$ [34]. By applying Thm. 5 to the iterates generated by Alg. 1 we have the following result.
@@
With the hypothes of Prop. 4 let ${\mathcal{H}}=W^{2,s}(\mathcal{X})$ with $s>d/2$. Let ${T}_{n}\in{HS}({\mathcal{F}},{\mathcal{G}})$ and $r_{n}\in{\mathcal{G}}$ characterized as in Prop. 3. Let $\pi_{0}(a|\cdot)\propto e^{\eta q_{0}(\cdot,a)}$ for $q_{0}$ such that $q_{0}(\cdot,a)\in{\mathcal{H}}$ any $a\in\mathcal{A}$. Then, for any $t\in\mathbb{N}$ the PMD iterates $\pi_{t}$ generated by Alg. 1 are such that $\pi_{t}(a|\cdot)\in{\mathcal{H}}$ and hence are $({\mathcal{G}},{\mathcal{F}})$-compatible.
The above corollary guarantees us that if we are able to learn our estimates for the action-value function in ${\mathcal{H}}$ a suitably regular Sobolev space, then POWR is well-defined. This is a necessary condition to then being able to study it’s theoretical behavior in our main result. We report the proofs of Thm. 5 and Cor. 6 in Sec. C.1.
@@
Thm. 7 shows that inexact PMD can behave comparably to its exact version provided that $\hat{q}_{\pi_{t}}$ are estimated with increasing accuracy, provided that the sequence of policies is $({\mathcal{G}},{\mathcal{F}})$-compatible, for example in the Sobolev-based setting of Cor. 6. Specifically, if $\left\|{q_{\pi_{t}}-\hat{q}_{\pi_{t}}}\right\|_{\infty}\leq O(1/t)$ for any $t\in\mathbb{N}$, the convergence rate of inexact PMD is of order $O(\log{T}/T)$, only a logarithmic factor slower than exact PMD. This means that we do not necessarily need a good approximation of the world model from the beginning but rather a strategy to improve upon such approximation as we perform more PMD iterations. This suggests adopting an alternating strategy between exploration (world model learning) and exploitation (PMD steps), as suggested in Sec. 4. We do not investigate this question in this work.
@@
Thm. 7 highlights the importance of studying the error of the estimator for the action-value functions produced by Alg. 1. These objects are obtained via the formula described in Prop. 3 in terms of ${T}_{n}$, $r_{n}$ and ${P}_{\pi}$. The exact $q_{\pi}$ has an analogous closed-form characterization in terms of ${T}$, $r$ and ${P}_{\pi}$ as expressed in 7 and motivating our operator-based approach. The following result compares these quantities in terms of the approximation errors of the world model and the reward function.
@@
Let ${T}_{n}$ and $r_{n}$ the empirical estimators of the transfer operator ${T}$ and reward function $r$ as defined in Prop. 3, respectively. Let also $\gamma\bigl{\|}{{T}_{n}}\bigr{\|}<\gamma^{\prime}<1$. Then,
$\bigl{\|}{\hat{q}_{\pi}-q_{\pi}}\bigr{\|}_{\infty}\leq\frac{1}{1-\gamma^{\prime}}\left[C_{\psi}\bigl{\|}{r_{n}-r}\bigr{\|}_{{\mathcal{G}}}+\frac{\gamma\bigl{\|}{r}\bigr{\|}_{\infty}}{1-\gamma}\bigl{\|}{{T}|_{{\mathcal{F}}}-{T}_{n}}\bigr{\|}_{{HS}}\right]$
In the result above, when applied to a function in ${\mathcal{G}}$, such as $r_{n}$, the uniform norm is to be interpreted as the uniform norm of the evaluation of such function, namely $\left\|{r_{n}}\right\|_{\infty}=\sup_{(x,a)\in\Omega}|\left\langle{r_{n}},{\psi(x,a)}\right\rangle|$, and analogously for ${T}_{n}$. The proof, reported in Sec. C.2, follows by first decomposing the difference $q_{\pi}-\hat{q}_{\pi}$ with the simulation lemma [26, Lemma 2.2] and then applying the triangular inequality for the uniform norm.
@@
We are ready to state the convergence result for Alg. 1. We consider the setting where the dataset used to learn ${T}_{n}$ (and $r_{n}$) is made of i.i.d. triplets $(x_{i},a_{i},x_{i}^{\prime})$ with $(x_{i},a_{i})$ sampled from a distribution $\rho\in{\mathcal{M}}_{1}^{+}(\Omega)$ supported on all $\Omega$ (such as the state occupancy measure (see e.g. [11] or Sec. A.3) of the uniform policy $\pi(\cdot|x)=1/|\mathcal{A}|$) and $x_{i}^{\prime}$ sampled from $\tau(\cdot|x_{i},a_{i})$. To guarantee bounds in uniform norm, the result makes a further regularity assumption, of the transfer operator (and the reward function)
@@
Let $\rho\in{\mathcal{M}}_{1}^{+}(\Omega)$ and $C_{\rho}$ the covariance operator $\sum_{a\in\mathcal{A}}\int_{\mathcal{X}}\rho(dx,a)\psi(x,a)\otimes\psi(x,a)$. The transition operator ${T}$ and the reward function $r$ are such that ${T}|_{\mathcal{F}}\in{HS}({\mathcal{F}},{\mathcal{G}})$ and $r\in{\mathcal{G}}$. Further, $\left\|{(T|_{\mathcal{F}})^{*}C_{\rho}^{-\beta}}\right\|_{{HS}}<\infty$ and $\left\|{C_{\rho}^{-\beta}r}\right\|_{{\mathcal{G}}}<\infty$ for some $\beta>0$.
Assumption 2 imposes a strong requirement to the so-called source condition, a quantity that describes how well the target objective of the learning process (here ${T}$ and $r$) “interact” with he sampling distribution. The assumption is always satisfied when the hypothesis space is finite dimensional (e.g. in the tabular RL setting) and imposes additional smoothness on ${T}$ and $r$ when belong to a Sobolev space. Equipped with this assumption, we can now state the convergence theorem for Alg. 1.
@@
Let $(\pi_{t})_{t\in\mathbb{N}}$ be a sequence of policies generated by Alg. 1 in the same setting of Cor. 6. If the action-value functions $\hat{q}_{\pi_{t}}$ are estimated from a dataset $(x_{i},a_{i};x_{i}^{\prime})_{i=1}^{n}$ with $(x_{i},a_{i})\sim\rho\in{\mathcal{M}}_{1}^{+}(\Omega)$ such that Asm. 2 holds with parameter $\beta$, the iterates of Alg. 1 converge to the optimal policy as
@@
The proof of Thm. 9 is reported in Appendix C and combines the results discussed in this section with fast convergence rates for the least-squares [35] and CME [18] estimators. In particular we fisrt use Thm. 5 to guarantee that the policies produced by Alg. 1 are all $({\mathcal{G}},{\mathcal{F}})$-compatible and therefore that applying Prop. 3 to obtain an estimator for the action-value function is well-defined. Then, we use Lemma 8 to study the approximation error of these estimators in terms of our estimates for the world model and the reward function. Bounds on these quantities are then used in the result for inexact PMD convergence in Thm. 7. We note here that since the latter results require convergence in uniform norm, we cannot leverage standard results for least-squares and CME convergence, which characterize convergence in $L_{2}(\Omega,\mu)$ and would only require Asm. 1 (Linear MDP) to hold. Rather, we need to impose Asm. 2 to guarantee faster rates in uniform norm.
@@ 6 Experimental results
We empirically evaluated POWR on classical Gym environments [36], ranging from discrete (FrozenLake-v1, Taxi-v3) to dense state spaces (MountainCar-v0). To ensure balancing between exploration and exploitation of our method, we alternated between running the environment with the current policy to collect samples for world model learning and running Alg. 1 for a number of steps to generate a new policy. Appendix D provides implementation details regarding this process as well as details on the kernel and hyperparameters used in both tabular and infinite-states settings.
Fig. 1 compares our approach with the performance of well-established baselines including A2C [37, 38], DQN [4, 38], TRPO [7, 38], and PPO [6, 38]. The figure reports the average cumulative reward obtained by the models on test environments with respect to the number of interactions with the MDP (timesteps in log scale in the figure) across 7 different training runs. In all plots, the horizontal dashed line represents the “success” threshold for the corresponding environment, according to official guidelines. We observe that our method outperforms all competitors by a significant margin in terms of sample complexity (i.e. reward achieved wrt number of timesteps executed) and, in the case of the Taxi-v3 environment, it avoids converging to a local optimum, in contrast to other methods such as A2C and TRPO. On the downside, we note that our method exhibits less stability than other approaches, particularly during the initial stages of the training process. This is arguably due to a sub-optimal interplay between exploration and exploitation, which will be the subject of future work.
@@ 7 Conclusions and Future Work
Our analysis opens avenues for further exploration. Firstly, extending PMD to infinite action spaces remains a challenge. While we introduced the operatorial perspective on RL for infinite state space settings, the PMD update with KL divergence requires approximation methods (e.g., Monte Carlo) whose impact on convergence requires investigation. Secondly, scalability to large environments requires adopting approximated yet efficient CME estimators like Nystrom [30] or reduced-rank regressors [39]. Thirdly, a question we touched upon only empirically, is whether alternating world model learning with inexact PMD updates benefits the exploration-exploitation trade-off. Studying this strategy’s impact on convergence is a promising future direction. Finally, a crucial question is generalizing our policy compatibility results beyond Sobolev spaces. Ideally, a representation learning process would identify suitable feature maps that guarantee compatibility with the PMD-generated policies while allowing for added flexibility in learning the world model.
@@ A.1 Auxiliary Lemma
The result is obvious if $A$ is invertible. More generally, we consider the following two application of the Sherman-Woodbury [40] formula
@@
Multiplying the two equation by $A$ respectively to the right and to the left, we obtain the desired result. ∎
@@ A.2 Markov operators and their properties
We recall here the notion of Markov operators, which is central for a number of results in the following. We refer to [25, Chapter 19] for more details on the topic.
@@
$\begin{split}&\textbf{a. }f\geq 0\in B_{b}(\mathcal{X})\implies{P}f\geq 0\in B_{b}(\mathcal{Y}),\\ &\textbf{b. }{P}\mathbf{1}_{\mathcal{X}}=\mathbf{1}_{\mathcal{Y}},\\ \end{split}$
@@
We recall that Markov operators are a convex subset of $\mathcal{L}(B_{b}(\mathcal{X}),B_{b}(\mathcal{Y}))$. Here we denote this space as $\mathcal{L}_{{\rm M}}(B_{b}(\mathcal{X}),B_{b}(\mathcal{Y}))$. Direct inspection of 5 and 6 shows that the transfer operator ${T}$ associated to an MDP and the policy operator ${P}_{\pi}$ associated to a policy $\pi$ are both Markov operators.
@@
In 6 we defined the policy operator ${P}_{\pi}$ associated to a policy $\pi$. It turns out that th e converse is also true, namely that any such Markov operator is a policy operator.
@@
Let ${P}\in\mathcal{L}_{{\rm M}}(B_{b}(\mathcal{X}),B_{b}(\Omega))$ be a Markov operator. Then there exists $\pi_{P}$, such that the associated policy operator corresponds to ${P}$, namely ${P}_{\pi_{P}}={P}$.
@@
Define the map $\pi_{P}:\mathcal{X}\to\mathcal{M}(\mathcal{A})$ taking value in the space of bounded Borel measures over $\mathcal{A}$ such that, for any $x\in\mathcal{X}$ and any $\mathcal{B}\subseteq\mathcal{A}$ Borel measurable subset
$\pi_{P}(\mathcal{B}|x)=({P}1_{\mathcal{X}\times\mathcal{B}})(x).$ (A.3)
We need to guarantee that for every $x\in\mathcal{X}$ the function $\pi_{P}(\cdot|x)$ is a signed measure. To show this, first note that the operation defined by $\pi_{P}$ is well-defined, since for any measurable set $\mathcal{B}$ the function $\mathbf{1}_{\mathcal{X}\times\mathcal{B}}$ is also measurable, making $\pi_{P}(\mathcal{B}|x)$ well defined as well. Moreover, since $\mathbf{1}_{\emptyset}(a)=0$ for any $a\in\mathcal{A}$, it implies that $\mathbf{1}_{\mathcal{X}\times\emptyset}=0$ and therefore $\pi_{P}(\emptyset|x)=0$ for any $x\in{\mathcal{X}}$. Finally, $\sigma$-additivity follows from the definition of indicator functions, namely $\mathbf{1}_{\bigcup_{i=1}^{\infty}\mathcal{B}_{i}}=\sum_{i=1}^{\infty}\mathbf{1}_{\mathcal{B}_{i}}$ for any family of pair-wise disjoint sets $(\mathcal{B}_{i})_{i=1}^{n}$, which implies $\pi_{P}\left(\bigcup_{i=1}^{\infty}\mathcal{B}_{i}|x\right)=\sum_{i=1}^{\infty}\pi_{P}(\mathcal{B}_{i}|x)$ for any $x\in\mathcal{X}$.
We now apply the two properties of Markov operators to show that $\pi_{P}$ takes values in ${\mathcal{M}}_{1}^{+}(\mathcal{A})$, namely it is a non-negative measure that sums to $1$. Since Markov operators map non-negative functions in non-negative functions and since $\mathbf{1}_{\mathcal{X}\times\mathcal{B}}\geq 0$ for any $\mathcal{B}\subseteq\mathcal{X}$, we have $\pi(\cdot|x)\geq 0$ as well for any $x\in\mathcal{X}$. Moreover, since $\Omega=\mathcal{X}\times\mathcal{A}$ and ${P}\mathbf{1}_{\Omega}=\mathbf{1}_{\mathcal{X}}$, we have
@@
for any $x\in\mathcal{X}$. Therefore $\pi_{P}(\cdot|x)$ is a probability measure for any $x\in\mathcal{X}$. Direct application of 6 shows that the associated policy operator corresponds to ${P}$, namely ${P}_{\pi_{P}}={P}$ as desired. ∎
Given the correspondence between policies and their Markov operator according to 6 and Proposition A.2, in the following we will denote the policy operator associated to a policy $\pi$ only ${P}$ where clear from context.
With the definition of Markov operator in place, we can now prove the following result introduced in the main paper.

### v2

@@ 5 Theoretical Analysis
We now show that POWR converges to the global maximizer of the RL problem in 1. To this end, we first identify a family of function spaces guaranteed to be compatible with the policies generated by Alg. 1. Then, we provide an extension of the result in [12] for inexact PMD to infinite state spaces $\mathcal{X}$, showing the impact of the action-value approximation error on the convergence rates. For the estimator introduced in Prop. 3, we relate this error to the approximation errors of $\mathsf{T}_{n}$ and $r_{n}$ leveraging the Simulation Lemma A.6. Finally, we use recent advancements in the characterization of CMEs’ fast learning rates to bound the sample complexity of these latter quantities, yielding error bounds for POWR.
@@
Let $\mathcal{X}\subset{\mathbb{R}}^{d}$ be a compact set and let ${\mathcal{H}}=W^{2,s}(\mathcal{X})$ be the Sobolev space of smoothness $s>0$ (see e.g. [37]). Let $\pi_{t}(a|\cdot)$ and $\hat{q}_{\pi_{t}}(\cdot,a)$ belong to ${\mathcal{H}}$ for any $a\in\mathcal{A}$ and $\pi_{t}(a|x)>0$ for any $x\in\mathcal{X}$. Then the policy $\pi_{t+1}$ solution to the PMD update in 14 belongs to ${\mathcal{H}}$.
According to Thm. 5, Sobolev spaces offer a viable choice for compatibility with PMD-generated policies. This observation is further supported by the fact that Sobolev spaces of smoothness $s>d/2$ are so-called reproducing kernel Hilbert spaces (rkhs) (see e.g. [38, Ch. 10]). We recall that rkhs are always naturally equipped with a $\phi:\mathcal{X}\to{\mathcal{H}}$ such that the inner product $\left\langle{\phi(x)},{\phi(x^{\prime})}\right\rangle=k(x,x^{\prime})$ defines a so-called reproducing kernel, namely a positive definite function that is (usually) efficient to evaluate, even if $\phi(x)$ is high or infinite dimensional. For example, ${\mathcal{H}}=W^{2,s}(\mathcal{X})$ with $s=\lceil\frac{d}{2}\rceil$ has associated kernel $k(x,x^{\prime})=e^{-\|x-x^{\prime}\|/\sigma}$ with bandwidth $\sigma>0$ [38]. By applying Thm. 5 to the iterates generated by Alg. 1 we have the following result.
@@
With the hypothesis of Prop. 4 let ${\mathcal{H}}=W^{2,s}(\mathcal{X})$ with $s>d/2$. Let $\mathsf{T}_{n}\in\mathsf{HS}({\mathcal{F}},{\mathcal{G}})$ and $r_{n}\in{\mathcal{G}}$ characterized as in Prop. 3. Let $\pi_{0}(a|\cdot)\propto e^{\eta q_{0}(\cdot,a)}$ for $q_{0}$ such that $q_{0}(\cdot,a)\in{\mathcal{H}}$ any $a\in\mathcal{A}$. Then, for any $t\in{\mathbb{N}}$ the PMD iterates $\pi_{t}$ generated by Alg. 1 are such that $\pi_{t}(a|\cdot)\in{\mathcal{H}}$ and hence are $({\mathcal{G}},{\mathcal{F}})$-compatible.
The above corollary guarantees us that if we are able to learn our estimates for the action-value function in a suitably regular Sobolev space ${\mathcal{H}}$, then POWR  is well-defined. This is a necessary condition to then being able to study its theoretical behavior in our main result. We report the proofs of Thm. 5 and Cor. 6 in Sec. C.1.
@@
Thm. 7 shows that inexact PMD can behave comparably to its exact version provided that 1) the action value functions $\hat{q}_{\pi_{t}}$ are estimated with increasing accuracy, and 2) that the sequence of policies is $({\mathcal{G}},{\mathcal{F}})$-compatible, for example in the Sobolev-based setting of Cor. 6. Specifically, if $\left\|{q_{\pi_{t}}-\hat{q}_{\pi_{t}}}\right\|_{\infty}\leq O(1/t)$ for any $t\in{\mathbb{N}}$, the convergence rate of inexact PMD is of order $O(\log{T}/T)$, only a logarithmic factor slower than exact PMD. This means that we do not necessarily need a good approximation of the world model from the beginning but rather a strategy to improve upon such approximation as we perform more PMD iterations. This suggests adopting an alternating strategy between exploration (world model learning) and exploitation (PMD steps), as suggested in Sec. 4. We do not investigate this question in this work.
@@
Thm. 7 highlights the importance of studying the error of the estimator for the action-value functions produced by Alg. 1. These objects are obtained via the formula described in Prop. 3 in terms of $\mathsf{T}_{n}$, $r_{n}$ and $\mathsf{P}_{\pi}$. The exact $q_{\pi}$ has an analogous closed-form characterization in terms of $\mathsf{T}$, $r$ and $\mathsf{P}_{\pi}$ as expressed in 7, and motivating our operator-based approach. The following result compares these quantities in terms of the approximation errors of the world model and the reward function.
@@
Let $\mathsf{T}_{n}$ and $r_{n}$ the empirical estimators of the transfer operator $\mathsf{T}$ and reward function $r$ as defined in Prop. 3, respectively. If $\mathsf{T}$ satisfies Asm. 1, $r\in{\mathcal{G}}$, and $\gamma\bigl{\|}{\mathsf{T}_{n}}\bigr{\|}<\gamma^{\prime}<1$, then, for every $({\mathcal{G}},{\mathcal{F}})$-compatible policy $\pi$
$\bigl{\|}{\hat{q}_{\pi}-q_{\pi}}\bigr{\|}_{\infty}\leq\frac{1}{1-\gamma^{\prime}}\left[\textrm{const}_{\psi}\bigl{\|}{r_{n}-r}\bigr{\|}_{{\mathcal{G}}}+\frac{\gamma\bigl{\|}{r}\bigr{\|}_{\infty}}{1-\gamma}\bigl{\|}{\mathsf{T}|_{{\mathcal{F}}}-\mathsf{T}_{n}}\bigr{\|}_{\mathsf{HS}}\right].$
In the result above, when applied to a function in ${\mathcal{G}}$, such as $r_{n}$, the uniform norm is to be interpreted as the uniform norm of the evaluation of such function, namely $\left\|{r_{n}}\right\|_{\infty}=\sup_{(x,a)\in\Omega}|\left\langle{r_{n}},{\psi(x,a)}\right\rangle|$, and analogously for $\mathsf{T}_{n}$. The proof, reported in Sec. C.2, follows by first decomposing the difference $q_{\pi}-\hat{q}_{\pi}$ with the simulation lemma [30, Lemma 2.2] and then applying the triangular inequality for the uniform norm.
@@
We are ready to state the convergence result for Alg. 1. We consider the setting where the dataset used to learn $\mathsf{T}_{n}$ (and $r_{n}$) is made of i.i.d. triplets $(x_{i},a_{i},x_{i}^{\prime})$ with $(x_{i},a_{i})$ sampled from a distribution $\rho\in\mathcal{P}(\Omega)$ supported on all $\Omega$ (such as the state occupancy measure (see e.g. [11] or Sec. A.3) of the uniform policy $\pi(\cdot|x)=1/|\mathcal{A}|$) and $x_{i}^{\prime}$ sampled from $\tau(\cdot|x_{i},a_{i})$. To guarantee bounds in uniform norm, the result makes a further regularity assumption, of the transfer operator (and the reward function)
@@
Let $\rho\in\mathcal{P}(\Omega)$ and $C_{\rho}$ the covariance operator
$C_{\rho}=\sum_{a\in\mathcal{A}}\int_{\mathcal{X}}\rho(dx,a)\psi(x,a)\otimes\psi(x,a).$
The transition operator $\mathsf{T}$ and the reward function $r$ are such that $\mathsf{T}|_{\mathcal{F}}\in\mathsf{HS}({\mathcal{F}},{\mathcal{G}})$ and $r\in{\mathcal{G}}$. Further, $\left\|{(T|_{\mathcal{F}})^{*}C_{\rho}^{-\beta}}\right\|_{\mathsf{HS}}<\infty$ and $\left\|{C_{\rho}^{-\beta}r}\right\|_{{\mathcal{G}}}<\infty$ for some $\beta>0$.
Asm. 2 imposes a strong requirement to the so-called source condition, a quantity that describes how well the target objective of the learning process (here $\mathsf{T}$ and $r$) “interact” with the sampling distribution. The assumption is always satisfied when the hypothesis space is finite dimensional (e.g. in the tabular RL setting) and imposes additional smoothness on $\mathsf{T}$ and $r$ when belonging to a Sobolev space. Equipped with this assumption, we can now state the convergence theorem for Alg. 1.
@@
Let $(\pi_{t})_{t\in{\mathbb{N}}}$ be a sequence of policies generated by Alg. 1 in the same setting of Cor. 6. If the action-value functions $\hat{q}_{\pi_{t}}$ are estimated from a dataset $(x_{i},a_{i};x_{i}^{\prime})_{i=1}^{n}$ with $(x_{i},a_{i})\sim\rho\in\mathcal{P}(\Omega)$ such that Asm. 2 holds with parameter $\beta$, the iterates of Alg. 1 converge to the optimal policy as
@@
The proof of Thm. 9 is reported in Appendix C and combines the results discussed in this section with fast convergence rates for the least-squares [39] and CME [21] estimators. In particular we first use Thm. 5 to guarantee that the policies produced by Alg. 1 are all $({\mathcal{G}},{\mathcal{F}})$-compatible and therefore that applying Prop. 3 to obtain an estimator for the action-value function is well-defined. Then, we use Lemma 8 to study the approximation error of these estimators in terms of our estimates for the world model and the reward function. Bounds on these quantities are then used in the result for inexact PMD convergence in Thm. 7. We note here that since the latter results require convergence in uniform norm, we cannot leverage standard results for least-squares and CME convergence, which characterize convergence in $L_{2}(\Omega,\mu)$ and would only require Asm. 1 (Linear MDP) to hold. Rather, we need to impose Asm. 2 to guarantee faster rates in uniform norm.
@@ 6 Experimental results
We empirically evaluated POWR on classical Gym environments [25], ranging from discrete (FrozenLake-v1,
Taxi-v3) to continuous state spaces (MountainCar-v0). To ensure balancing between exploration and exploitation of our method, we alternated between running the environment with the current policy to collect samples for world model learning and running Alg. 1 for a number of steps to generate a new policy. Appendix D provides implementation details regarding this process as well as additional results.
Fig. 1 compares our approach with the performance of well-established baselines including A2C [40], DQN [4], TRPO [7], and PPO [6]. The figure reports the average cumulative reward obtained by the models on test environments with respect to the number of interactions with the MDP (timesteps in log scale in the figure) across 7 different training runs. In all plots, the horizontal dashed line represents the “success” threshold for the corresponding environment, according to official guidelines. We observe that our method outperforms all competitors by a significant margin in terms of sample complexity, that is, the reward achieved after a given number of timesteps. In the case of the Taxi-v3 environment, it avoids converging to a local optimum, in contrast every other method with the exception of DQN. On the downside, we note that our method exhibits less stability than other approaches, particularly during the initial stages of the training process. This is arguably due to a sub-optimal interplay between exploration and exploitation, which will be the subject of future work.
@@ 7 Conclusions and Future Work
Our analysis opens avenues for further exploration. Firstly, extending PMD to infinite action spaces remains a challenge. While we introduced the operatorial perspective on RL for infinite state space settings, the PMD update with KL divergence requires approximation methods (e.g., Monte Carlo) whose impact on convergence requires investigation. Secondly, scalability to large environments requires adopting approximated yet efficient CME estimators like Nyström [34] or reduced-rank regressors [41, 42]. Thirdly, a question we touched upon only empirically, is whether alternating world model learning with inexact PMD updates benefits the exploration-exploitation trade-off. Studying this strategy’s impact on convergence is a promising future direction. Finally, a crucial question is generalizing our policy compatibility results beyond Sobolev spaces. Ideally, a representation learning process would identify suitable feature maps that guarantee compatibility with the PMD-generated policies while allowing for added flexibility in learning the world model.
Acknowledgments
We acknowledge financial support from NextGenerationEU and MUR PNRR project PE0000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)”, EU grant ELISE (GA no 951847) and EU Project ELIAS (GA no 101120237).
@@ A.1 Auxiliary Lemma
The result is obvious if $A$ is invertible. More generally, we consider the following two applications of the Sherman-Woodbury [43] formula
@@
Multiplying the two equations by $A$ respectively to the right and to the left, we obtain the desired result. ∎
@@ A.2 Markov operators and their properties
We recall here the notion of Markov operators, which is central for a number of results in the following. We refer to [29, Chapter 19] for more details on the topic.
@@
$\begin{split}&\textbf{a. }f\geq 0\in B_{b}(\mathcal{X})\implies\mathsf{P}f\geq 0\in B_{b}(\mathcal{Y}),\\ &\textbf{b. }\mathsf{P}\mathbf{1}_{\mathcal{X}}=\mathbf{1}_{\mathcal{Y}},\\ \end{split}$
@@
We recall that Markov operators are a convex subset of $\mathcal{L}(B_{b}(\mathcal{X}),B_{b}(\mathcal{Y}))$. Here we denote this space as $\mathcal{L}_{{\rm M}}(B_{b}(\mathcal{X}),B_{b}(\mathcal{Y}))$. Direct inspection of 5 and 6 shows that the transfer operator $\mathsf{T}$ associated to an MDP and the policy operator $\mathsf{P}_{\pi}$ associated to a policy $\pi$ are both Markov operators.
@@
In 6 we defined the policy operator $\mathsf{P}_{\pi}$ associated to a policy $\pi$. It turns out that the converse is also true, namely that any such Markov operator is a policy operator.
@@
Let $\mathsf{P}\in\mathcal{L}_{{\rm M}}(B_{b}(\mathcal{X}),B_{b}(\Omega))$ be a Markov operator. Then there exists $\pi_{\mathsf{P}}$, such that the associated policy operator corresponds to $\mathsf{P}$, namely $\mathsf{P}_{\pi_{\mathsf{P}}}=\mathsf{P}$.
@@
Define the map $\pi_{\mathsf{P}}:\mathcal{X}\to\mathcal{M}(\mathcal{A})$ taking value in the space of bounded Borel measures over $\mathcal{A}$ such that, for any $x\in\mathcal{X}$ and any $\mathcal{B}\subseteq\mathcal{A}$ Borel measurable subset
$\pi_{\mathsf{P}}(\mathcal{B}|x)=(\mathsf{P}1_{\mathcal{X}\times\mathcal{B}})(x).$ (A.3)
We need to guarantee that for every $x\in\mathcal{X}$ the function $\pi_{\mathsf{P}}(\cdot|x)$ is a signed measure. To show this, first note that the operation defined by $\pi_{\mathsf{P}}$ is well-defined, since for any measurable set $\mathcal{B}$ the function $\mathbf{1}_{\mathcal{X}\times\mathcal{B}}$ is also measurable, making $\pi_{\mathsf{P}}(\mathcal{B}|x)$ well defined as well. Moreover, since $\mathbf{1}_{\emptyset}(a)=0$ for any $a\in\mathcal{A}$, it implies that $\mathbf{1}_{\mathcal{X}\times\emptyset}=0$ and therefore $\pi_{\mathsf{P}}(\emptyset|x)=0$ for any $x\in{\mathcal{X}}$. Finally, $\sigma$-additivity follows from the definition of indicator functions, namely $\mathbf{1}_{\bigcup_{i=1}^{\infty}\mathcal{B}_{i}}=\sum_{i=1}^{\infty}\mathbf{1}_{\mathcal{B}_{i}}$ for any family of pair-wise disjoint sets $(\mathcal{B}_{i})_{i=1}^{n}$, which implies $\pi_{\mathsf{P}}\left(\bigcup_{i=1}^{\infty}\mathcal{B}_{i}|x\right)=\sum_{i=1}^{\infty}\pi_{\mathsf{P}}(\mathcal{B}_{i}|x)$ for any $x\in\mathcal{X}$.
We now apply the two properties of Markov operators to show that $\pi_{\mathsf{P}}$ takes values in $\mathcal{P}(\mathcal{A})$, namely it is a non-negative measure that sums to $1$. Since Markov operators map non-negative functions in non-negative functions and since $\mathbf{1}_{\mathcal{X}\times\mathcal{B}}\geq 0$ for any $\mathcal{B}\subseteq\mathcal{X}$, we have $\pi(\cdot|x)\geq 0$ as well for any $x\in\mathcal{X}$. Moreover, since $\Omega=\mathcal{X}\times\mathcal{A}$ and $\mathsf{P}\mathbf{1}_{\Omega}=\mathbf{1}_{\mathcal{X}}$, we have
@@
for any $x\in\mathcal{X}$. Therefore $\pi_{\mathsf{P}}(\cdot|x)$ is a probability measure for any $x\in\mathcal{X}$. Direct application of 6 shows that the associated policy operator corresponds to $\mathsf{P}$, namely $\mathsf{P}_{\pi_{\mathsf{P}}}=\mathsf{P}$ as desired. ∎
Given the correspondence between policies and their Markov operator according to 6 and Proposition A.2, in the following we will denote the policy operator associated to a policy $\pi$ only $\mathsf{P}$ where clear from context.
With the definition of the Markov operator in place, we can now prove the following result introduced in the main paper.